# Unc-TTP: A Method for Classifying LLM Uncertainty to Improve In-Context Example Selection

### v2

@@ Abstract
Nowadays, Large Language Models (LLMs) have demonstrated exceptional performance across various downstream tasks. However, it is challenging for users to discern whether the responses are generated with certainty or are fabricated to meet user expectations. Estimating the uncertainty of LLMs is particularly challenging due to their vast scale and the lack of white-box access. In this work, we propose a novel Uncertainty Tripartite Testing Paradigm (Unc-TTP) to classify LLM uncertainty, via evaluating the consistency of LLM outputs when incorporating label interference into the sampling-based approach. Based on Unc-TTP outputs, we aggregate instances into certain and uncertain categories. Further, we conduct a detailed analysis of the uncertainty properties of LLMs and show Unc-TTP’s superiority over the existing sampling-based methods. In addition, we leverage the obtained uncertainty information to guide in-context example selection, demonstrating that Unc-TTP obviously outperforms retrieval-based and sampling-based approaches in selecting more informative examples. Our work paves a new way to classify the uncertainty of both open- and closed-source LLMs, and introduces a practical approach to exploit this uncertainty to improve LLMs performance.
Unc-TTP: A Method for Classifying LLM Uncertainty
to Improve In-Context Example Selection
@@ 1 Introduction
"Real knowledge is to know the extent of one’s ignorance." —— Confucius In the rapidly evolving field of artificial intelligence (AI), the quest for generic AI assistants to solve problems with truthful and reliable information remains challenging. Although Large Language Models (LLMs) are capable of fulfilling user requests for various downstream tasks, they struggle with admitting their knowledge boundary (Xu et al., 2024a). Additionally, they sometimes fabricate statements that are difficult to discern from falsehoods, posing trustworthiness issues (Maynez et al., 2020; Alkaissi and McFarlane, 2023; Ji et al., 2023). These challenges push us to consider two critical questions: (1) How can we determine whether the LLM is certain about its answer? (2) Can we leverage the uncertainty property of the model to improve its performance?
Much effort has been dedicated to understanding when an LLM is certain of its responses. One line of work employs white-box access to the internal model information to provide interpretability for deep neural networks (Gal and Ghahramani, 2016; Amini et al., 2020; Kuhn et al., 2023). These methods become increasingly complex and resource-intensive as models scale to LLMs (Doshi-Velez and Kim, 2017; Krishna et al., 2023), and are unsuitable for closed-source LLMs like GPT-3.5 and GPT-4 without the involvement of a white-box proxy model.
To broadly evaluate different LLMs (both open- and close-sourced), another line of work turns to black-box methods by analyzing the input-output behavior of models to infer their internal reasoning process. For instance, some research has employed prompt engineering methods to assess whether LLMs have truly grasped the essence of knowledge (Wang et al., 2023a; Turpin et al., 2023a; Xie et al., 2024; Wei et al., 2024a). These studies suggest that LLMs exhibit a sign of sycophancy, tending to waver between answers when exposed to interference. While this behavior highlights the presence of uncertainty, existing studies have merely observed it without leveraging it as an indicator to measure the LLM’s uncertainty. Further research has explored sampling-based methods to evaluate LLM certainty Wang et al. (2023b); Cole et al. (2023); Manakul et al. (2023b), which assesses the answer consistency between multiple decoding paths. However, we argue that this approach measures uncertainty too loosely. In this paper, we incorporate label injection into the sampling-based method and classify the uncertainty from the consistency perspective, offering a more fine-grained classification.
Several attempts have been made to exploit the uncertainty (or certainty) properties of LLMs to improve their performance. Xu et al. (2024a); Xie et al. (2024) collect data where the model’s performance meets reliability expectations, and then train the model for optimizing reliability. However, these methods rely heavily on the annotated data and demand significant computational resources. Mo et al. (2024) introduce an example selection approach to few-shot information extraction by incorporating both correct and incorrect demonstrations, leveraging self-consistency Wang et al. (2023b) to refine example selection. However, given the loose criteria of sampling-based self-consistency, we believe that a stricter classification of instances could further enhance in-context learning (ICL).
In this paper, we propose a novel Uncertainty Tripartite Testing Paradigm (Unc-TTP), by enumerating all the possibilities of the LLM under triple-testing scenarios {no-label, right-label, wrong-label} and classifying the model uncertainty in terms of response consistency. Specifically, we obtain the original outputs of LLMs as well as the output after injecting the right or wrong labels as interference. The triplet assigns each instance into one of $82^{3}=$ 8 categories since the model can either answer correctly or incorrectly in each of the three settings, respectively. On top of these 8 categories, we consider the rest to be uncertain instances of the LLM, except for the instances where the model is unwaveringly right and wrong. We conduct a detailed analysis of LLM uncertainty classified by Unc-TTP and the sampling-based method such as self-consistency on three subjective datasets. Our proposed Unc-TTP reveals the inherent wavering nature of the LLM in the face of external disturbances, but it also exhibits varying degrees of resistance depending on the dataset and capability of the model. Compare to the sampling-based method, our Unc-TTP provides a much stricter classification of LLM certainty.
Inspired by the uncertainty sampling strategy in active learning (Lewis and Gale, 1994; Lewis and Catlett, 1994; Settles, 2009), we employ uncertainty to enhance LLMs performance via ICL. Specifically, we utilize the uncertainty categories identified by Unc-TTP to guide the selection of in-context examples. Compared to prior research on example selection, which typically relies on time-consuming retrieval-based methods to select the optimal examples for each test instance (Robertson and Zaragoza, 2009; Liu et al., 2021; Rubin et al., 2022; Zhang et al., 2022), our approach offers a one-example-for-all solution. In our method, example selection is completed in a single round, with the selected examples applicable to all test instances. This approach significantly boosts inference efficiency and delivers notable performance gains compared to the retrieval-based method BM25 (Robertson and Zaragoza, 2009), which requires carefully selecting examples for each individual test instance. It also surpasses the random and sampling-based method by a large margin.
@@
  • We propose a novel paradigm, Uncertainty Tripartite Testing Paradigm (Unc-TTP), to test the model uncertainty for both open- and closed-source LLMs, offering a new perspective on classifying uncertainty.
  • We conduct a detailed analysis of the instance distribution categorized as “certain" and “uncertain" by both the sampling-based method and Unc-TTP across various LLMs. Our experiments reveal that both methods effectively pinpoint areas of uncertainty within the LLMs, while stronger LLMs exhibit greater certainty. Additionally, when LLMs encounter disagreement with users, they tend to choose the neutral option to align with user expectations.
  • Based on Unc-TTP, we propose an uncertainty-based in-context example selection strategy that enhances LLM’s performance in the few-shot setting. Compared to traditional retrieval-based approaches, our self-guided, one-example-for-all solution yields better performance in an efficient way. Additionally, the uncertainty criterion of our Unc-TTP is much stricter than that of the sampling-based method, further enhancing ICL performance.
@@
2 Related Work
2.1 The Manifestations of LLM Uncertainty
While LLMs have performed well in various downstream tasks, their reliability is still questioned, particularly in highly specialized domains such as medicine  (Alkaissi and McFarlane, 2023; Shen et al., 2023). Previous work has identified sycophancy and hallucinations as two main reasons for the unreliability of LLMs, which can be attributed to the manifestations of model uncertainty.
Sycophancy
The sycophancy of LLMs is that they frequently wrongly admit mistakes and give biased feedback when questioned by the user  (Sharma et al., 2023). By using LLMs for debates (Wang et al., 2023a), introducing misleading elements into prompts (Turpin et al., 2023b, a; Xie et al., 2024; Wei et al., 2024a), or injecting the answer before engaging in Chain-of-Thought (CoT) reasoning (Xu et al., 2024b), these findings collectively suggest that sycophancy may be a prevalent trait among LLMs, indicating their inability to maintain a consistent stance.
The patterns of sycophancy have been consistently observed across various LLMs and targeted for reduction in previous works (Sharma et al., 2023; Yang et al., 2023; Wei et al., 2024a, b). To take a step forward, we recognize the wavering behavior behind sycophancy as an indicator of uncertainty within LLMs.
Hallucination
LLMs often generate content that conflicts with sources or lacks factual verification, known as hallucinations (Li et al., 2023a; Huang et al., 2023). Xu et al. (2024a) argue that it occurs when LLMs attempt to answer questions beyond their knowledge boundaries. To address this, they train a reward model to enable the LLM to refuse to answer questions about which it is uncertain. Similarly, Xiong et al. (2024) attempt to get LLMs to verbalize their confidence in their answers but reveal a tendency towards overconfidence. We attribute these issues to the LLMs’ inability to express their uncertainty without additional assistance. Nevertheless, we can still unveil their uncertainty through their extrinsic manifestations.
2.2 ICL and Example Selection Strategy
ICL is a paradigm that allows LLMs to learn tasks given only a few examples as demonstrations without updating model parameters (Dong et al., 2023; Brown et al., 2020).
The most critical factor affecting a model’s ICL performance is the demonstration examples (Kumar and Talukdar, 2021; Lu et al., 2022). Significant efforts have been made to select the best in-context examples for LLMs. Liu et al. (2021) retrieve in-context examples based on their semantic similarity to test samples. Nguyen and Wong (2023) suggest selecting examples by calculating their influence. Qin et al. (2023) iteratively select diverse yet strongly correlated examples as demonstrations. Robertson and Zaragoza (2009) employ BM25 to select ICL examples based on their relevance scores. However, these methods select in-context examples for each test instance dynamically, which can be resource-intensive.
Some works utilize examples of what LLMs have done wrong to introduce more information into demonstrations. Mo et al. (2024) leverage both correct and incorrect samples, with incorrect samples selected through self-consistency Wang et al. (2023b), to enhance the diversity of ICL demonstrations. Krishna et al. (2023) select examples through a weak proxy model, and amplify LLM performance by leveraging these difficult examples and their post hoc explanation in ICL. However, these methods either involve sampling-based approaches, which we consider less precise for identifying LLM blind spots or rely on proxy models with white-box access. In contrast, we propose a self-guided, consistency-based method by taking advantage of LLM uncertainty.
3 Classifying LLM Uncertainty
Despite the sycophantic tendencies of LLMs when confronted with users’ skepticism or misinformation, they still exhibit varying degrees of resistance in their decision-making processes. Therefore, we propose an Uncertainty Tripartite Testing Paradigm (Unc-TTP), a method designed to assess the uncertainty of LLMs, which can be applied to both open- and closed-source models. The schematic of our proposed paradigm is illustrated in Figure 1.
3.1 The Uncertainty Tripartite Testing Paradigm
@@
Unc-TTP involves asking LLMs to answer a question under three independent settings: (1) providing with no label, (2) providing with the right label, and (3) providing with the wrong label, where the labels are sourced from benchmark datasets. We denote this tripartite paradigm as {no-label, right-label, wrong-label} for the sake of brevity.
Under the no-label setting, we instruct the LLM with “Your job is to determine whether the text is A or B", where “A" and “B" are the placeholder for the actual labels in the dataset. We thereby get the model answer that is derived entirely from the LLM’s own knowledge and reasoning ability, without any external guiding interference.
Under the right-label and wrong-label settings, we provide LLM with either a right or wrong label, intending to lure the model into making the decision we are steering it towards. These settings expose areas where the model is not confident and wavers between answers. However, recent work has shown that LLMs can be easily misled into believing in falsehoods and providing answers that align with the user’s misguided intentions (Wang et al., 2023a). Therefore, slavishly guiding the model by concatenating the externally injected correct (or incorrect) label with the task prompt does not necessarily reflect the true uncertainty of the model but rather an inclination toward sycophancy. To diminish this possibility, we prompt the model with “Your job is to determine whether the text is A or B by reference to the given label, which presents the ground truth of the text as A or B. Despite having to refer to the provided labels, you should still have your own thinking and do not change your stance so easily".
3.2 Uncertainty Category Labeling
@@
After testing all three settings, we obtain three results corresponding to {no-label, right-label, wrong-label}, each with two possible outcomes: the model answering correctly or incorrectly. Marking the instances that the model answers correctly as 1 and incorrectly as 0, we get a total of eight possible outcomes, denoted as {000, 001, 010, 011, 100, 101, 110, 111}. These eight categories can be further grouped into two types:
@@
Certain When the model provides consistent answers for an instance under all three settings, we consider the model is certain about the instance. Category 000 (denoted as CertainW), and 111 (denoted as CertainR), fall within this type, where the model outputs unwaveringly wrong in the former and unwaveringly right in the latter.
@@
Uncertain In the rest of the cases where the model has wavering answers, including 001, 110, 011, 100, 101, and 110, we consider the model to be uncertain about the instance. Taking 011 as an example, it represents a data instance where the model answers incorrectly with the no-label setting and correctly with the right-label and wrong-label settings.
4 Assessing LLM Uncertainty
4.1 Experimental Setup
Model Configurations
We perform our experiments with GPT-3.5 and GPT-4  (OpenAI, 2024), llama-2-7b-chat (Touvron et al., 2023) and mistral-7b-instruct (Jiang et al., 2023), where all generations are done via greedy decoding.
Datasets
We select three text classification tasks that involve subjective decision-making for our experiments, according to the subjectivity ranking Li et al. (2023b). Sarcasm Headlines (SH) (Misra and Arora, 2023), a high-quality news headlines sarcasm detection dataset collected from two online news websites. Humor Speech (HS) (Annamoradnejad and Zoghi, 2020), a dataset consisting of short texts with both humorous and non-humorous content. Financial Phrasebank (FP) (Malo et al., 2014), a sentiment classification task that divides finance-related sentences as positive, negative, or neutral.
Comparison Method
Following Qin et al. (2024), we apply self-consistency Wang et al. (2023b) with 3 decoding paths (temperature 0.7) as sampling-based method. We label instances where all three sampling results are right as CertainR. We additionally label instances where all three samplings are wrong as CertainW, and those who sometimes get one or two right as Uncertain, which aligns with the uncertainty classification of Unc-TTP.
Implementation Details
To ensure fairness, we conduct this part of the experiment using the training set, while remaining blind to the validation and test sets. We adopt Accuracy as the evaluation metric. In the wrong-label setting of Unc-TTP, for datasets with more than two labels, such as FP, we randomly select an incorrect label to serve as the provided label.
4.2 Results and Analysis
Unc-TTP Uncertainty Performance
Figure 2: Evaluation results of Unc-TTP for four models.
Figure 3: Distribution of certain and uncertain instances on the three datasets, labelled with Unc-TTP (the top line) and the sampling-based method (the bottom line). The certain ones, unwaveringly right, are represented by Instances 111 and unwaveringly wrong by Instances 000. The uncertain ones, wavering are represented by the rest of the categories.
The evaluation results of Unc-TTP on three benchmarks are shown in Figure 3. LLM shows more compliance with correct labels. For the setting that provides labels, all models except Llama-2 on the SH dataset show gullibility to the given labels. Particularly, in the right-label setting, some models nearly reach 100% agreement with the given label. However, for the wrong-label setting, the model exhibits varying levels of resistance to enticement. Besides, the more capable the model is, the better it maintains accuracy in the wrong-label setting, although it does not converge to the accuracy achieved in the no-label setting. This indicates that unless the LLM is very certain of its answer, it reacts quite prudently when the labels provided deviate considerably from its own answers, as adopting the user’s point of view seems to be a safer option.
LLMs tend to make compromises when conservative options are offered, choosing the neutral option to accommodate the user’s preference. In the FP dataset, all models exhibit relatively steady performance under the wrong-label setting. A possible explanation for this observation is that FP is a classification dataset with three labels, where the distinctions between neutral and positive or negative are subtle. LLMs inherently tend to favor the more conservative neutral answers. When their initial beliefs are challenged in the wrong-label setting, they are inclined to prefer the neutral option, potentially aligning with user preferences. We observe that most LLMs exhibit a tendency to produce conservative answers in the no-label setting. This tendency becomes pronounced in the wrong-label setting, where the number of neutral answers doubles, further substantiating our conclusion.
Uncertainty Properties of LLMs
To further investigate the degree of uncertainty across different LLMs, we present the distribution of certain and uncertain instances for each model in Figure 3.
Stronger LLMs exhibit greater certainty. Intuitively, the more robust the model, the less likely it is to waver between answers. In Unc-TTP, despite external interference, GPT-4 maintains its stance to a large extent, whereas the other models exhibit resilience towards interference but are not as pronounced. A similar trend is observed in the sampling-based method, where stronger models exhibit a lower tendency to waver. We attribute this to the model’s confidence in its answers. Stronger models possess better logical reasoning and greater certainty, allowing them to maintain their stance regardless of label injection or repetitive sampling.
Label interference improves the strictness of classifying certainty standards. Compared to the sampling-based method, Unc-TTP injects right/wrong label into the prompt, making it much more difficult for LLMs to persist in its answer given how much it wavers in Figure 3. Therefore, the standard of Unc-TTP’s certainty is stricter than that of the sampling-based method. It is worth noting that we are not disapproving of the certainty delineated by the sampling-based approach; it is just that such a loose criterion might lead to an inadequate exposition of uncertain instances.
Uncertainty Category Distribution
We visualize the respective quantities of instances for each category in Figure 4. The distribution reveals significant unevenness, indicating that each LLM and dataset possesses a unique uncertainty property. Interestingly, the number of instances in the categories that do not match human intuition, such as 001, 100, and 101, is notably low, as expected. For instance, if a model misclassifies an instance both without guidance and with correct guidance, it should have a very low likelihood of answering correctly when provided with incorrect labels.
Figure 4: Number of examples corresponding to the respective labels for each model on the three datasets. The number of instances per category is distributed unevenly.
5 Applying Uncertainty to Example Selection
In this section, we provide an insight into the real-world applications of Unc-TTP through ICL. Inspired by the uncertainty sampling strategy in active learning, where models improve their learning efficiency by actively selecting informative data for training, we creatively select instances from the uncertainty categories identified by Unc-TTP to serve as ICL demonstrations, aiming to verify the impact of Unc-TTP-guided uncertainty on LLMs.
5.1 Experiment Settings
Example Selection Details

### v4

# Unlocking the Power of LLM Uncertainty for Active In-Context Example Selection

@@ Abstract
Large Language Models (LLMs) have shown remarkable performance across a wide range of downstream tasks. However, it is challenging for users to discern whether the responses of LLM are generated with certainty or are fabricated to meet user expectations. In this paper, we introduce Uncertainty Tripartite Testing Paradigm (Unc-TTP), a novel method for classifying LLM uncertainty by leveraging output inconsistency. Specifically, Unc-TTP performs three rounds of sampling under varying label injection interference, enumerating all possible outcomes, and uses the degree of output inconsistency as the indicator of the LLM’s intrinsic uncertainty. To validate the effectiveness of this inconsistency-defined uncertainty, we draw inspiration from Active Learning, comparing the informativeness of actively selected in-context examples. Our experiments show that uncertainty examples selected via Unc-TTP are more informative than certainty examples. Furthermore, the Unc-TTP-guided uncertainty-based active example selection strategy outperforms existing methods, highlighting its effectiveness in classifying LLM uncertainty and enhancing in-context learning. This work not only underscores the potential of inconsistency-based uncertainty classification for both open- and closed-source LLMs but also presents a practical approach for leveraging uncertainty to improve LLM performance in real-world tasks.
Unlocking the Power of LLM Uncertainty for
Active In-Context Example Selection
@@ 1 Introduction
"Real knowledge is to know the extent of one’s ignorance." —— Confucius
@@
In the rapidly evolving field of artificial intelligence (AI), developing generic AI assistants that provide truthful and reliable information remains a challenge. Although Large Language Models (LLMs) are capable of fulfilling user requests for various downstream tasks, they struggle with admitting their knowledge boundary (Xu et al., 2024a). Additionally, they sometimes fabricate statements that are difficult to discern from falsehoods, causing trustworthiness issues (Maynez et al., 2020; Alkaissi and McFarlane, 2023; Ji et al., 2023), especially when confronted with users’ skepticism or misinformation (Sharma et al., 2023). These challenges push us to consider three critical questions: (1) How can we assess the uncertainty of an LLM’s response through its external behavior? (2) How can we validate the effectiveness of the assessed uncertainty? (3) Can we utilize the assessed uncertainty to enhance the performance of LLMs?
Significant efforts have been devoted to investigating the external input-output behavior of LLM to assess their underlying reasoning abilities and conjecture whether LLMs have truly grasped the essence of knowledge (Wang et al., 2023a; Turpin et al., 2023a; Xie et al., 2024; Wei et al., 2024a). These studies indicate that LLMs display some tendencies of sycophancy, often wavering between answers when confronted with human interference. While current works are still trying to find ways to eliminate this sycophantic behavior, our approach takes a different direction. We capitalize on this sycophancy to develop a method for classifying uncertainty, leveraging the identified uncertainty to enhance LLM performance.
We begin with a key observation from a simple experiment in which we sample LLM responses three times for the exact same instance, under slightly different conditions—specifically, injecting no (w/o), right, or wrong label into the prompt (Section 3.1). Surprisingly, we found that although LLMs exhibit a clear tendency of sycophancy when faced with label injection, they still show varying degrees of resistance to this behavior. In some cases, LLMs consistently maintain their stance across all three samplings, with stronger models holding their answers more firmly. Building on this observation, we hypothesize that the LLM output inconsistency can serve as an indicator of its intrinsic uncertainty.
Accordingly, we propose a fine-grained Uncertainty Tripartite Testing Paradigm (Unc-TTP), which enumerates all possible outcomes of the LLM under triple-testing scenarios {no-label, right-label, wrong-label} and classifies LLM uncertainty levels based on the response consistency. Consequently, the Unc-TTP triplet assigns each instance to one of $82^{3}=$ 8 categories, as the LLM can either provide a correct or incorrect answer in each of the three settings. Given these eight categories, we classify any instance, where the LLM answers correct or incorrect waveringly, to be uncertain.
Basically, our proposed Unc-TTP is a kind of inconsistency-defined uncertainty. Drawing inspiration from the uncertainty sampling strategy in Active Learning (AL)  (Lewis and Gale, 1994; Lewis and Catlett, 1994; Settles, 2009), we apply Unc-TTP’s classified uncertainty for active in-context example selection in ICL experiments. The preliminary study demonstrates that inconsistency-defined uncertainty examples are more informative than certainty-based ones, providing strong evidence for its potential for boosting LLM performance.
Furthermore, based on three subjective text classification tasks, we conduct experiments by applying other existing strategies for active in-context example selection. Our Unc-TTP-selected uncertainty examples significantly outperform the previous best strategy, Similarity Margatina et al. (2023), with average accuracy improvements of 3.7%, 1.2%, and 1.9% for Llama-2, Mistral, and GPT-3.5, respectively, showcasing the effectiveness of our approach in enhancing ICL performance. We also explore the scalability and transferability of Unc-TTP classified uncertainty in selecting in-context examples, demonstrating its robustness and providing compelling evidence that uncertainty defined by Unc-TTP, which is calculated based on the LLM output inconsistency, really reveals the intrinsic uncertainty of LLMs.
@@
  • To the best of our knowledge, we are the first to experimentally demonstrate that output inconsistency can act as an indicator of the LLMs’ intrinsic uncertainty via the perspective of active learning.
@@
  • We propose a novel paradigm, Uncertainty Tripartite Testing Paradigm (Unc-TTP), to test the LLM uncertainty for both open- and closed-source LLMs, offering a new insight on classifying uncertainty.
  • Taking advantage of Unc-TTP, we propose an uncertainty-based active in-context example selection strategy, which yields greater performance gains than previous approaches in ICL.
2 Related Work
2.1 Uncertainty Behavior of LLM
While LLMs have performed well across a variety of downstream tasks, their reliability remains a topic of concern, especially in highly specialized domains such as medicine (Alkaissi and McFarlane, 2023; Shen et al., 2023). This is particularly evident when LLMs are questioned by the user and they frequently wrongly admit mistakes and give biased feedback (Sharma et al., 2023). The phenomenon where a model seeks human approval in undesirable or inappropriate ways is referred to as sycophancy (Perez et al., 2022).
Recent studies have highlighted the prevalence of sycophantic behavior in LLMs, with evidence found in various scenarios such as debates (Wang et al., 2023a), prompt manipulation to introduce misleading elements (Turpin et al., 2023a, b; Xie et al., 2024; Wei et al., 2024a), and preemptively injecting answers before engaging in Chain-of-Thought (CoT) reasoning (Xu et al., 2024b). These findings collectively suggest that sycophancy may be a prevalent trait among LLMs, indicating their inability to disregard user preference and maintain a consistent stance.
While others attempt to mitigate the sycophancy of LLM (Sharma et al., 2023; Yang et al., 2023; Wei et al., 2024a, b), our research proposes that the wavering behavior behind sycophancy can reflect the intrinsic uncertainty of LLMs. This uncertainty, in turn, could serve as a useful indicator for performing active in-context example selection.
2.2 ICL and Example Selection Strategy
ICL is a paradigm that allows LLMs to learn tasks given only a few examples as demonstrations without updating model parameters (Brown et al., 2020; Dong et al., 2023). In ICL, the quality of the demonstration examples is the most critical factor affecting LLM performance (Kumar and Talukdar, 2021; Lu et al., 2022).
Research on selecting optimal in-context examples for LLMs has been an ongoing area of study. Margatina et al. (2023) draws inspiration from the most common AL strategies, including uncertainty sampling (Gonen et al., 2024), diversity sampling (Yu et al., 2023), and similarity-based selection (Liu et al., 2021), to identify the most informative examples. However, these approaches fail to account for the unique characteristics of LLMs, particularly the intrinsic uncertainty revealed through sycophantic behavior. As we demonstrate in Section 5.1, this wavering behavior provides a more effective indicator for selecting in-context examples compared to traditional AL methods.
Other works have explored the use of incorrect examples, aiming to introduce more diverse information into the demonstrations. For instance, Mo et al. (2024) leverages incorrect examples selected through self-consistency (Wang et al., 2023b) to enhance diversity in examples. Krishna et al. (2023) selects incorrect examples using a weak proxy model and generates post-hoc explanations for these difficult cases. However, these methods either overlook the value of inconsistency—an important aspect that can help identify LLM blind spots—or depend on proxy models with white-box access.
In this paper, we propose a self-guided, uncertainty-based in-context example selection method that outperforms existing active ICL baselines.
3 Preliminary and Methodology
Despite the sycophantic tendencies of LLMs when confronted with users’ skepticism or misinformation, can they disregard users’ potentially bias information and maintain their stance? To investigate this question, we start with a simple test, whose observations lead to the development of a new methodology of classifying LLM uncertainty to improve in-context example selection: the Uncertainty Tripartite Testing Paradigm (Unc-TTP).
3.1 A Motivating Observation
The simple test is conducted under three settings: We prompt LLMs with regular query, query with right label and query with wrong label, using the prompt shown in Table 1, the LLMs and datasets described in Sections 4.1 and 4.2. The accuracy of four models is shown in Figure 2. Surprisingly, we observed varying degrees of resistance to sycophantic behavior in LLMs during their decision-making processes. In some cases, even when provided with right or wrong labels interference, the LLMs maintained their stance unwaveringly, revealing that, in some particular instances, LLMs can firmly maintain their stance across all three settings. This behavior might reflect their intrinsic uncertainty and certainty.
To further investigate this potential uncertainty (define by output inconsistency) across different LLMs, we present the distribution of wavering and unwavering instances for each model, selected using the settings described above (top), alongside the results from the vanilla method, which samples outputs three times under the no label injection setting using temperature sampling (bottom), as shown in Figure 2.
Derived from the results of Figure 2, we make the following observations. Stronger LLMs exhibit greater output consistency, with a larger percentage of them being able to unwaveringly maintain their answers across three samplings. We attribute this to the LLM’s confidence in its answers. Stronger models possess better logical reasoning ability, allowing them to maintain their stance regardless of label injection or temperature sampling. Moreover, compared to the vanilla sampling-based method, injecting right/wrong label into the prompt, making it much more difficult for LLMs to persist in its answer, therefore improves the strictness of identifying the actual instances where LLMs are certain.
Building on these key observations, we hypothesize that the LLM output inconsistency can serve as an indicator of its uncertainty. In order to prove the validity of this hypothesis, we draw inspiration from uncertainty sampling in AL and adopt AL principle for in-context example selection, where we select the most informative instances with our inconsistency-defined uncertainty as selection strategy. Since LLMs exhibit varying behaviors for different injected labels (see Appendix A), this underscores the need for a fine-grained uncertainty classification method. To address this, we propose Unc-TTP.
Figure 1: Accuracy of four models with and without label injection. See Appendix A for a detailed analysis.
(a) gpt-4
w/ label injection
(b) gpt-3.5
w/ label injection
(c) llama-2
w/ label injection
(d) mistral
w/ label injection
(e) gpt-4-vanilla
(f) gpt-3.5-vanilla
(g) llama-2-vanilla
(h) mistral-vanilla
Figure 2: Distribution of instances where the LLM remains unwavering (certain) and wavering (uncertain) across three rounds of sampling, with {no (w/o), right, wrong} label injection (top) and without label injection, using only vanilla temperature sampling (bottom), across three datasets.
Prompt
no label setting (w/o label injection) Your job is to determine whether the text is {A} or {B}.
right / wrong label setting
Your job is to determine whether the text is {A} or {B} by reference to the given label, which presents the ground truth of the text as {A} or {B}. Despite having to refer to the provided labels, you should still have your own thinking and do not change your stance so easily.
Table 1: Prompts used in Unc-TTP, where {A} and {B} denote the actual label from the training data.
3.2 The Uncertainty Tripartite Testing Paradigm
@@
The schematic of our proposed Unc-TTP is illustrated in Figure 3. It involves asking LLMs to answer a question under three independent settings: (1) providing with no (w/o) label, (2) providing with the right label, and (3) providing with the wrong label, where the labels are sourced from benchmark datasets. We denote this tripartite paradigm as {no-label, right-label, wrong-label} for the sake of brevity.
At the same time, to encourage the model to think independently, we prompt the model to keep its own stances when facing label interference, and the specific prompt is shown in Table 1.
The rationale behind no-label setting is that we want to get the model answer that is derived entirely from the LLM’s own knowledge and reasoning ability, without any external guiding interference. As for the right-label and wrong-label settings, we provide LLM with either a right or wrong label, intending to lure the model into making the decision we are steering it towards. These settings expose areas where the model is not certain.
Uncertainty Category Labeling
@@
After testing all three settings, for each instance, we obtain three results corresponding to {no-label, right-label, wrong-label}, each with two possible outcomes: the model answering correctly or incorrectly. Marking the instances that the model answers correctly as 1 and incorrectly as 0, we get a total of eight possible outcomes, denoted as {000, 001, 010, 011, 100, 101, 110, 111}. These eight categories can be further grouped into two types:
@@
Certain: When the model provides consistent answers for an instance under all three settings, we consider the model to be certain with respect to the instance. The categories 000 (denoted as CerW), and 111 (denoted as CerR) fall into this type, where the model’s outputs are unwaveringly wrong in the former and unwaveringly right in the latter.
@@
Uncertain: In the cases where the model provides wavering answers, including 001, 110, 011, 100, 101, and 110, we consider the model to be uncertain about the instance. Taking 011 as an example, it represents a data instance where the model answers incorrectly with the no-label setting and correctly with the right-label and wrong-label settings. We denote these instances as Uncertain (Unc).
3.3 Example Selection with Unc-TTP
Taking advantage of Unc-TTP, we assign each instance in the training set with a specific uncertainty categories, and the category distribution is shown in Figure 6. The distribution reveals significant unevenness, indicating that each LLM and dataset possesses a unique uncertainty property. In this section, we select examples based on the fine-grained uncertainty categories classified by Unc-TTP.